# Unified Multimodal Interleaved Document Representation for Retrieval

### v2

@@ Abstract
Jaewoo Lee1,2∗   Joonho Ko1∗   Jinheon Baek1∗   Soyeong Jeong1   Sung Ju Hwang1,2 KAIST1 DeepAuto2 jwlee8877@gmail.com, {joonho.ko, jinheon.baek, starsuzi}@kaist.ac.kr
@@ 1 Introduction
Information Retrieval (IR) is the task of fetching relevant documents from a large corpus in response to an input query, which plays a critical role in various real-world applications including web search engines and question-answering systems (Shah et al., 2019; Lewis et al., 2020; Guu et al., 2020). Over the years, IR methods have evolved significantly, with approaches broadly categorized into sparse and dense retrieval paradigms. Specifically, sparse retrieval methods Robertson et al. (1994); Jones (2004) focus on lexical overlap between queries and documents; meanwhile, dense retrieval methods Karpukhin et al. (2020); Xiong et al. (2021) utilize neural embeddings to represent queries and documents in a continuous vector space. Note that, recently, dense retrieval methods have gained more popularity over sparse methods due to their capability to capture semantic nuances and context beyond simple keyword matching, leading to multiple successes with improved performance.
@@
Despite their huge successes, existing (dense) retrieval methods face a couple of severe challenges. First, they primarily rely on the textual data for document embedding and retrieval, overlooking the fact that modern documents often contain multimodal content, such as images and tables (beyond the plain text), which can carry critical information that may be essential for accurately understanding and retrieving the relevant documents. To be specific, a diagram within a medical article can more effectively represent the structure of a molecule or the progression of a disease, offering more clarity that would be difficult to achieve with text alone, and omitting such multimodal content can lead to an incomplete understanding (and potentially inaccurate retrieval) of the documents. Also, the segmentation of long documents into discrete passages, which is commonly employed by existing retrieval models to handle the length limitation for embeddings, may prevent models from capturing the full context and the intricate relationships between different parts of the document, ultimately leading to suboptimal retrieval performance. It is worthwhile noting that, concurrent to our work, while there has been recent work that screen captures the document and then embed its screenshots (to consider different modalities in a unified format) (Faysse et al., 2024; Ma et al., 2024), not only its content (such as paragraphs, images, and tables) can be fragmented into different sub-images, leading to the loss of contextual coherence across the entire document, but also the visual representation of text may hinder the model’s ability to capture the semantic relationships present in the original textual data, while increasing the image resolution leads to the concern on the memory requirements.
@@
In this work, we introduce a novel approach to holistically represent documents for IR, which addresses the aforementioned challenges by representing and retrieving the documents interleaved with multiple modalities in a unified manner (See Figure 1). Specifically, our method revolves around the recent advance of Vision-Language Models (VLMs), which enable the processing and integration of multimodal content (such as text, images, and tables) directly into a single token sequence, thereby preserving the context and relationships between various parts of the document, unlike the previous approaches that rely on the fragmented visual representations. Also, in cases where the number of tokens in a document is large and exceeds the capacity of a single context window of VLMs, we propose a strategy to segment the document into passages, each represented within the token limit, and combine these passage embeddings into a unified document representation. This strategy differs from existing IR approaches that independently represent and retrieve at the passage level, potentially losing the overall document context. Lastly, to accurately identify only the relevant sections within the retrieved lengthy document, we introduce a reranking mechanism that is trained to pinpoint the passage most pertinent to the query (among all the other passages within the document), allowing for both the coarse-grained document-level matching and fine-grained passage-level retrieval. We refer to our overall framework as Interleaved Document Information Retrieval System (IDentIfy).
@@
We experimentally validate the effectiveness of IDentIfy on four benchmark datasets, considering both the text-only and multimodal queries. On a battery of tests conducted, we observe that our approach substantially outperforms relevant baselines that consider only the uni-modality for document representations, thanks to the consideration of multimodal content. Further, we find that the strategy to represent the whole document with its single representation (by merging embeddings of its splits) is superior to the approach of individually representing them for document retrieval, but also performing reranking over the sections of the retrieved document is superior to the approach of directly retrieving those sections, which confirm the efficacy of the proposed retrieval and reranking pipeline for document and passage retrieval, respectively.
@@ 2 Related Work
Information Retrieval (IR) involves finding documents relevant to a query, which plays a crucial role in applications such as search and question-answering (Zhu et al., 2023; Gao et al., 2023; Ram et al., 2023; Shi et al., 2024; Jeong et al., 2024a). Earlier IR approaches measured the similarity between queries and documents based on their lexical term matching, such as BM25 and TF-IDF (Robertson et al., 1994; Jones, 2004). Yet, these methods often struggled to capture the semantic nuances beyond surface-level term overlaps. To overcome this, along with advancements in language models (Devlin et al., 2019; Liu et al., 2019), there have been dense retrieval approaches that embed both the queries and documents into a shared dense vector space (Karpukhin et al., 2020; Xiong et al., 2021), enabling the calculation of semantic similarity between them more effectively by capturing the deeper contextual information. Yet, previous studies have mainly focused on enhancing the textual representations of queries and documents, while overlooking the multimodal nature of documents beyond text, which can potentially provide richer context and aid in more accurate retrieval (Liu et al., 2021; Jeong et al., 2024b).
@@
Recent studies in IR have expanded the focus from purely text-based retrieval models to those that consider other modalities, such as images (Radford et al., 2021; Xiao et al., 2024), tables (Herzig et al., 2021; Chen et al., 2024) and graphs (Baek et al., 2023); however, the majority of these approaches (Zhou et al., 2024; Long et al., 2024; Lerner et al., 2024; Nowak et al., 2024; Caffagni et al., 2024) have primarily explored how to process the multimodal queries, meanwhile, they often overlook the equally important multimodal characteristics of the documents being retrieved. In efforts to handle diverse multimodal elements within documents, there are concurrent studies that have proposed to capture screenshots of documents, such as PDFs (Faysse et al., 2024; Cho et al., 2024) or Wikipedia web pages (Ma et al., 2024), and subsequently encoding them through vision models (Ding et al., 2024). However, these methods are not only limited by factors, such as image resolution and computational memory, constraining their application to documents longer than a single page [^1] For instance, Ma et al. (2024) requires processing 9.8k image tokens just to process a single-page document, and it results in 2TB of storage for handling the entire Wikipedia corpus, which may not be practical., but also fall short by treating the diverse modalities within a document as a single visual entity, leading to suboptimal document representations that may fail to effectively capture the nuanced interdependence between text and images. Also, while there are concurrent studies Jiang et al. (2024b); Lin et al. (2024) that consider images and text as retrieval targets, they primarily focus on representing image-text pairs and their retrieval, rather than addressing the holistic representation of documents that include multiple images and another modality (tables). Finally, all the aforementioned work does not address the issue of splitting documents into smaller fragments (passages or sub-images), which may disrupt the holistic contextual view of the entire document.
@@
Recently developed Vision-Language Models (VLMs) have emerged as a powerful tool for jointly processing visual and textual data, which combine the image understanding capabilities of visual encoders (Radford et al., 2021; Zhai et al., 2023) with the advanced reasoning abilities of language models (OpenAI, 2022, 2023a). These models have achieved remarkable performance across diverse vision-language (VL) tasks (such as image captioning and visual question answering) (Dai et al., 2023; OpenAI, 2023b), with the substantially limited attention on their applications to IR. We note that the latest developments in this field have particularly focused on enabling VLMs to handle interleaved, multimodal content, which involves a mixed sequence of images and text (Zhang et al., 2023; Li et al., 2024b). In particular, LLaVA-NeXT-Interleave (Li et al., 2024b) introduces a fine-tuning approach that specifically enhances the VLMs’ capacity to understand complex interleavings of multiple images and text within a single context. Drawing inspiration from these advances, we propose to harness the capabilities of VLMs to create unified embeddings for documents interleaved with text and images (as well as tables).
@@ 3.1 Preliminaries
We begin with preliminaries, formally explaining information retrieval and vision-language models.
@@
IR is formally defined as the task of identifying a set of relevant documents $\{{\bm{d}}_{1},{\bm{d}}_{2},\ldots,{\bm{d}}_{k}\}\subseteq\mathcal{D}$ from a large corpus $\mathcal{D}$, given an input query ${\bm{q}}$. Here, each query ${\bm{q}}$ and document ${\bm{d}}$ are represented as a sequence of tokens: ${\bm{q}}=[q_{1},q_{2},\ldots,q_{n}]$ and ${\bm{d}}=[d_{1},d_{2},\ldots,d_{m}]$, and traditional IR approaches typically consider these tokens as purely textual elements. However, we propose to extend this assumption to have the tokens of both the textual and visual content, to capture the multimodal nature of many real-world documents. Then, this new extension raises important questions of how can both the textual and visual content be represented within a unified token framework, and how can these multimodal tokens be seamlessly integrated and encoded for document representations. To answer them, we harness the power of recent vision-language models below.
Figure 2: Overview of the proposed IDentIfy. (a): In our document retriever, a query encoder represents a query (purple), and sections are encoded with a section encoder whose embeddings averaged to form a document representation (blue). Contrastive learning loss (red) is used for training the document retriever. (b): Reranker scores query-section relevance with the concatenation of the query and section, trained using Binary Cross-Entropy loss.
@@
We now turn to describing Vision-Language Models (VLMs), which are designed to jointly encode the textual and visual information in a unified token framework. These models are generally comprised of two main components: a visual encoder and a language model, interconnected through a projection layer. Specifically, given the document that may contain interleaved modalities (e.g., text and images), the visual encoder extracts high-level visual features from (multiple) images embedded within the document, mapping them into a latent space. Then, these visual features are transformed into a sequence of visual tokens via the projection layer, represented as follows: $\mathbf{V}\!\in\!\mathbb{R}^{V\!\times\!d_{\text{emb}}}$, where $V$ denotes the visual token length and $d_{\text{emb}}$ is the token dimension size. Similarly, for the textual content embedded within the document, the language model uses a word embedding layer to convert the input text into a sequence of tokens, as follows: $\mathbf{L}\!\in\!\mathbb{R}^{L\!\times\!d_{\text{emb}}}$, where $L$ denotes the token length of text.
In this work, we also propose to account for tables that are an integral modality for holistically representing the full content of documents. However, in contrast to text and images that have dedicated processing layers within the VLM architectures, tables do not have a specific representation layer. Nevertheless, we argue that recent VLMs are pre-trained on diverse web data, and subsequently they are implicitly learned to handle the table structures formatted in HTML. Consequently, we treat HTML-format table data as a linearized sequence of HTML words, applying the same word embedding layer as is used for plain text. To be formal, this process converts the table content into table tokens, as follows: $\mathbf{T}\!\in\mathbb{R}^{T\!\times\!d_{\text{emb}}}$, where $T$ is the token length of the table. Lastly, once extracted, the visual tokens, text tokens, and table tokens are concatenated (to form a unified token sequence) and then passed through the remaining layers of VLMs, to capture both uni- and cross-modal relationships across different modalities, ultimately enabling the comprehensive understanding of the documents.
@@ 3.2 Retriever
We now explain how we design a retriever specifically tailored for multimodal interleaved document retrieval. In particular, our approach leverages a VLM capable of processing text, images, and tables within a single document. Further, following the standard practice of existing retrieval architectures (Karpukhin et al., 2020; Xiong et al., 2021), we use a dual-encoder structure, which consists of a query encoder and a section encoder, both are based on the VLM, illustrated in Figure 2 (a).
@@
The remaining step to discuss here is how to train those two query and section encoders for IR. Recall that the goal of the retriever is to assess a relevance score between the query and the document. To achieve this goal, we use a contrastive learning loss based upon the query and document representations, whose objective is to assign higher similarity scores to relevant documents (positive samples) and lower scores to irrelevant ones (negative samples) for the query, formulated as follows:
$\mathcal{L}_{\text{retriever}}\!$ $=\!-\frac{1}{B}\sum^{B}_{i=1}\log\left(\frac{\phi\left(\mathbf{Z}_{\mathrm{Q}_{i}},\mathbf{Z}_{\mathrm{D}_{i}}\right)}{\sum^{B}_{j=1}\phi\left(\mathbf{Z}_{\mathrm{Q}_{i}},\mathbf{Z}_{\mathrm{D}_{j}}\right)}\right),$
$\phi\left(\mathbf{a},\;\mathbf{b}\right)\!$ $=\!\exp\left(\frac{\mathbf{a}^{\top}\mathbf{b}}{\lVert\mathbf{a}\rVert\lVert\mathbf{b}\rVert}\right),$ (1)
where $B$ is the batch size during the training phase. Here, by minimizing $\mathcal{L}_{\text{retriever}}$, the retriever learns to optimize the similarity between queries and their relevant documents, enabling the retrieval of the most pertinent documents (among all) for the given input query during inference.
@@ 3.3 Reranker
To enable fine-grained retrieval within documents beyond the retrieval of documents themselves, we introduce a section-level reranking mechanism that identifies the section most relevant to the query. In particular, once the document is retrieved, the objective of the reranker $f_{\mathrm{R}}$ is to pinpoint the specific sections within the document that best match the query. We also note that this reranker is similarly operationalized with the VLM along with a binary classifier on top of it, which directly measures the relevance of each query-section pair (Figure 2 (b)).
@@
$\mathcal{L}_{\text{reranker}}$ $\!=\!\sum_{i=1}^{B}\sum_{j=1}^{S_{i}}\!\frac{1}{BS_{i}}\ell\left(\mathbf{y}_{{\bm{s}}_{i,j}},\;f_{\mathrm{R}}\left(\left[{\bm{q}},\;\hat{{\bm{s}}}_{i,j}\right]\right)\right),$
$\ell\left(y,\hat{y}\right)$ $\!=\!-\left[y\log\hat{y}\!+\!(1\!-\!y)\log(1\!-\!\hat{y})\right],$ (2)
@@ 4 Experiments
Table 1: Results with different document formats for retrieval.
Format R@1 R@10 R@100 MRR@10

### v3

@@ Abstract
Jaewoo Lee1∗   Joonho Ko2∗   Jinheon Baek2∗   Soyeong Jeong2   Sung Ju Hwang2,3 University of North Carolina Chapel Hill1   KAIST2   DeepAuto3 jwoolee@cs.unc.edu, {joonho.ko, jinheon.baek, sungju.hwang}@kaist.ac.kr
@@ 1 Introduction
Information Retrieval (IR) is the task of fetching relevant documents from a large corpus in response to a query, which plays a critical role in various real-world applications including web search engines and question-answering systems (Shah et al., 2019; Lewis et al., 2020; Guu et al., 2020). Over the years, IR methods have evolved significantly, broadly categorized into sparse and dense retrieval paradigms. Specifically, sparse retrieval methods Robertson et al. (1994); Jones (2004) focus on lexical overlap between queries and documents; meanwhile, dense retrieval methods Karpukhin et al. (2020); Xiong et al. (2021) utilize neural embeddings to represent queries and documents in a continuous vector space. Note that, recently, dense retrieval methods have gained more popularity over sparse methods due to their capability to capture semantic nuances and context beyond simple keyword matching.
@@
Despite their successes, existing (dense) retrieval methods face a couple of severe challenges. First, they primarily rely on the textual data for document embedding and retrieval, overlooking the fact that modern documents often contain multimodal content, including images and tables (beyond the plain text), which can carry information that may be essential for accurately understanding and retrieving the relevant documents Li et al. (2024c). For instance, a diagram within a medical article can more effectively represent the structure of a molecule or the progression of a disease, offering more clarity that would be difficult to achieve with text alone, and omitting such multimodal content can lead to an incomplete understanding (and potentially inaccurate retrieval) of the documents. Also, the segmentation of long documents into discrete passages, which is commonly employed by existing retrieval models to handle the length limitation for embeddings (Karpukhin et al., 2020; Xiong et al., 2021), may prevent models from capturing the full context and the intricate relationships between different parts of the document, ultimately leading to suboptimal retrieval performance (Dong et al., 2024; Jiang et al., 2024c). Notably, concurrent to our work, while there has been recent work that screen captures the document and then embed its screenshots (to consider different modalities in a unified format) (Faysse et al., 2024; Ma et al., 2024), not only its content (such as paragraphs, images, and tables) can be fragmented into different sub-images, leading to the loss of contextual coherence across the entire document, but also the visual representation of text may hinder the model’s ability to capture the semantic relationships present in the original textual data, and increasing image resolution raises concerns on memory requirements.
@@
To tackle these challenges, we introduce a novel approach to holistically represent documents for IR, representing and retrieving documents interleaved with multiple modalities in a unified manner (illustrated in Figure 1). Specifically, it revolves around the recent advance of Vision-Language Models (VLMs), which enable the processing and integration of multimodal content (such as text, images, and tables) directly into a single token sequence, thereby preserving the context and relationships between various parts of the document, unlike prior methods that rely on the fragmented visual representations. Additionally, in cases where the number of tokens in a document is large and exceeds the capacity of a single context window of VLMs, we propose a strategy to segment the document into passages, each represented within the token limit, and combine these passage embeddings into a unified document representation. This strategy differs from existing approaches that independently represent and retrieve at the passage level, potentially losing the overall document context. Lastly, to accurately identify only the relevant sections within the retrieved lengthy document, we introduce a reranking mechanism that is trained to pinpoint the passage most pertinent to the query (among all the other passages within the document), allowing for both the coarse-grained document-level matching and fine-grained passage-level retrieval. We refer to our overall framework as Interleaved Document Information Retrieval System (IDentIfy).
@@
We experimentally validate the effectiveness of IDentIfy on four benchmark datasets, considering both text-only and multimodal queries. On a battery of tests conducted, we observe that our approach substantially outperforms relevant baselines that consider only the uni-modality or certain facets of multi-modality, thanks to the holistic consideration of interleaved multimodal contents. Furthermore, we find that the strategy to represent the whole document with its single representation (by merging embeddings of its splits) is superior to the approach of individually representing them for document retrieval, but also performing reranking over the sections of the retrieved document is superior to the approach of directly retrieving those sections, confirming the efficacy of proposed coarse-to-fine retrieval and reranking pipeline for document and passage retrieval, respectively.
@@ 2 Related Work
IR involves finding documents relevant to a query, which plays a crucial role in applications such as search and question-answering (Zhu et al., 2023; Gao et al., 2023; Ram et al., 2023; Shi et al., 2024; Jeong et al., 2024b). Earlier IR approaches measured the similarity between queries and documents based on their lexical term matching, such as BM25 and TF-IDF (Robertson et al., 1994; Jones, 2004). However, these methods struggled to capture semantic nuances beyond surface-level term overlaps. Recently, along with advancements in language models (Devlin et al., 2019; Liu et al., 2019), there have been dense retrieval methods that embed both queries and documents into a shared dense vector space (Karpukhin et al., 2020; Xiong et al., 2021), enabling the calculation of semantic similarity between them more effectively by capturing the deeper contextual information. Yet, previous studies have mainly focused on enhancing the textual representations of queries and documents, while overlooking the multimodal nature of documents beyond text, which can provide richer context and aid in more accurate retrieval (Liu et al., 2021; Jeong et al., 2024a).
@@
Recent studies in IR have expanded the focus from purely text-based retrieval models to those that consider other modalities, such as images (Radford et al., 2021; Xiao et al., 2024), tables (Herzig et al., 2021; Chen et al., 2024) and graphs (Baek et al., 2023); however, the majority of these approaches (Zhou et al., 2024; Long et al., 2024; Lerner et al., 2024; Nowak et al., 2024; Caffagni et al., 2024) have primarily explored how to process the multimodal queries, and overlooked the equally important multimodal characteristics of the documents being retrieved. In efforts to handle diverse multimodal elements within documents, there are concurrent studies that have proposed to capture screenshots of documents, such as PDFs (Faysse et al., 2024; Cho et al., 2024) or Wikipedia web pages (Ma et al., 2024), and subsequently encoding them through vision models (Ding et al., 2024). Yet, these methods are not only limited by factors, such as image resolution and computational memory, constraining their application to documents longer than a single page [^1] It requires processing 9.8k image tokens just to process a single-page document, and it results in 2TB of storage for handling the entire Wikipedia corpus, which may not be practical., but also fall short by treating the diverse modalities within a document as a single visual entity, leading to document representations that may fail to effectively capture the nuanced interdependence between text and images. Also, while there are concurrent studies Jiang et al. (2024d); Lin et al. (2024) that consider images and text as retrieval targets, they primarily focus on representing image-text pairs and their retrieval, rather than addressing the holistic representation of documents that include multiple images and another modality (tables). Finally, all the aforementioned work does not address the issue of splitting documents into smaller fragments (passages or sub-images), which may disrupt the holistic contextual view of the entire document.
@@
Recently developed VLMs have emerged as a powerful tool for jointly processing visual and textual data, which combine the image understanding capabilities of visual encoders (Radford et al., 2021; Zhai et al., 2023) with the advanced reasoning abilities of language models (OpenAI, 2022, 2023a). They have achieved remarkable performance across diverse vision-language tasks (such as image captioning and visual question answering) (Dai et al., 2023; OpenAI, 2023b), with substantially limited attention on their applications to IR. We note that the latest developments in this field have particularly focused on enabling VLMs to handle interleaved, multimodal content, involving a mixed sequence of images and text (Zhang et al., 2023; Li et al., 2024b). In particular, LLaVA-NeXT-Interleave (Li et al., 2024b) introduces a fine-tuning approach that specifically enhances the VLMs’ capacity to understand complex interleavings of multiple images and text in a single context. Drawing inspiration from these advances, we propose to harness their capabilities to create unified embeddings for documents interleaved with text and images (and tables).
@@ 3.1 Preliminaries
We begin with formally explaining IR and VLMs.
@@
[…]
Figure 2: Overview of the proposed IDentIfy. (a): In our document retriever, a query encoder represents a query (purple), and sections are encoded with a section encoder whose embeddings are averaged to form a document representation (blue). Contrastive learning loss (red) is used for training the document retriever. (b): Reranker scores query-section relevance with the concatenation of the query and section, trained using Binary Cross-Entropy loss.
@@
To answer them, we now turn to describing VLMs, which are designed to jointly encode the textual and visual information in a unified token framework. These models are generally comprised of two main components: a visual encoder and a language model, interconnected through a projection layer. Specifically, given the document that may contain interleaved modalities (e.g., text and images), the visual encoder extracts high-level visual features from images embedded within the document, mapping them into a latent space. Then, these visual features are transformed into a sequence of visual tokens via the projection layer, represented as follows: $\mathbf{V}\!\in\!\mathbb{R}^{V\!\times\!d_{\text{emb}}}$, where $V$ denotes the visual token length and $d_{\text{emb}}$ is the token dimension size. Similarly, for the textual content embedded within the document, the language model uses a word embedding layer to convert the input text into a sequence of tokens, as follows: $\mathbf{L}\!\in\!\mathbb{R}^{L\!\times\!d_{\text{emb}}}$, where $L$ denotes the text token length.
In this work, we also propose to account for tables that are the integral modality to holistically represent the full content of documents. Yet, unlike text and images that have dedicated processing layers within VLM architectures, tables do not have a specific representation layer. Nevertheless, we argue that VLMs are pre-trained on diverse web data, and subsequently learned implicitly to handle the table structures formatted in HTML. Consequently, we treat HTML-format table data as a linearized sequence of HTML words, applying the same word embedding layer as is used for plain text. To be formal, this process converts the table content into table tokens, as follows: $\mathbf{T}\!\in\mathbb{R}^{T\!\times\!d_{\text{emb}}}$, where $T$ is the token length of the table. Lastly, once extracted, the visual tokens, text tokens, and table tokens are concatenated (into a unified token sequence) and then passed through the remaining layers of VLMs, to capture both uni- and cross-modal relationships across different modalities, ultimately enabling the comprehensive understanding of the documents.
@@ 3.2 Retriever
We now explain how we design a retriever specifically tailored for multimodal interleaved document retrieval. In particular, our approach leverages a VLM capable of processing text, images, and tables within a single document. Further, following the standard practice of existing retrieval architectures (Karpukhin et al., 2020; Xiong et al., 2021), we use a dual-encoder structure, which consists of a query encoder and document (or section) encoder, both are based on VLMs, illustrated in Figure˜2 (a).
@@
The remaining step is to train those two query and section encoders. Recall that the goal of the retriever is to assess a relevance score between the query and the document. To achieve this, we use a contrastive learning loss based upon the query and document representations, whose objective is to assign higher similarity scores to relevant documents (positive samples) and lower scores to irrelevant ones (negative samples) for the query, as follows:
[…]
where $B$ is the batch size. By minimizing $\mathcal{L}_{\text{retriever}}$, the retriever learns to optimize the similarity between queries and their relevant documents, enabling the retrieval of the most pertinent documents for the given input query during inference.
@@ 3.3 Reranker
To enable fine-grained retrieval within documents beyond the retrieval of documents themselves, we introduce a section-level reranking mechanism that identifies the section most relevant to the query. In particular, once the document is retrieved, the objective of the reranker $f_{\mathrm{R}}$ is to pinpoint the specific sections within the document that best match the query. We also note that this reranker is similarly operationalized with the VLM along with a binary classifier on top of it, which directly measures the relevance of each query-section pair (Figure˜2 (b)).
@@
[…]
@@ 4 Experiments
4.1 Experimental Setups
Datasets
We evaluate the proposed IDentIfy on four benchmark datasets designed for multimodal IR that require understanding of both textual and visual cues within queries and documents, as follows: Encyclopedic-VQA (Mensink et al., 2023) is a large-scale benchmark for multimodal Visual Question Answering (VQA) with queries linked to specific Wikipedia sections and includes both textual and multimodal queries; InfoSeek (Chen et al., 2023) is a knowledge-intensive VQA dataset with multimodal questions generated from Wikidata triples that include diverse entities such as landmarks, animals, and food; ViQuAE (Lerner et al., 2022) involves both textual and multimodal queries about human entities, linked to annotated Wikipedia sections, making it ideal for evaluating section reranking; Open-WikiTable (Kweon et al., 2023) targets open-domain table QA by identifying documents or sections containing relevant tables. We provide more details on datasets in Appendix˜A.
Baselines
To comprehensively validate IDentIfy, we compare it against two categories of baselines:
  • Conventional VLM Baselines: We consider earlier VLMs, which are not capable of jointly processing text and images, such as CLIP Radford et al. (2021) and BLIP (Li et al., 2022). Also, we consider the approaches, such as UniIR (Wei et al., 2024), which is built on top of them and fine-tuned with a contrastive loss (Equation˜1). These baselines serve as reference points to assess performance gains from recent VLM advances rather than serving as direct competitors.
  • Baselines with Different Document Representations: We further consider existing approaches, representing documents in various ways. Entity and Abstract baselines retrieve documents based on their titles and summaries, respectively, using high-level textual cues. Text-only baselines utilize the full textual content of documents for retrieval (Caffagni et al., 2024; Wang et al., 2024). Text & Table and Text & Image baselines leverage tables and first image of documents alongside the text, respectively (Jiang et al., 2024a; Lin et al., 2024; Jiang et al., 2024d). IDentIfy is our model that holistically represents multimodal content (text, images, and tables) in documents. All baselines share the same recent VLMs as IDentIfy, allowing for a controlled comparison focused on document representation strategies.
Method R@1 R@10 R@100 MRR@10
CLIP-VIT-L-14
Zero-Shot 1.9 6.3 13.9 3.1
UniIR + Text-Only 3.8 20.6 50.3 7.7
UniIR + Text & Image 5.8 21.5 48.5 10.0
BLIP-Large
Zero-Shot 0.0 0.0 0.0 0.0
UniIR + Text-Only 9.8 36.9 71.4 16.3
UniIR + Text & Image 9.9 23.9 60.7 13.5
LLaVA-NeXT-Interleave-0.5B